OpenStudy (anonymous):

Can someone help me differntiate cos pix/2

OpenStudy (anonymous):

\[\cos pix/2\]

OpenStudy (matt101):

d/dx(cos((pi/2)x)=(-pi/2)sin((pi/2)x)

OpenStudy (matt101):

I'm assuming that was the question...

OpenStudy (anonymous):

\[{ d \cos ({\pi \over 2}x) \over dx } = - \sin ({\pi \over 2}x) * {\pi \over 2}\]

OpenStudy (anonymous):

There was no parentheses in the orginal question, would what you wrote still be the correct differntiation?

OpenStudy (matt101):

yup